# Stability-Informed Initialization of Neural Ordinary Differential Equations

### v1

@@ 1 Introduction
Learning dynamic systems from observations has a long history of applications, particularly in time-series analysis and prediction. This research domain is intrinsically connected to classical system identification and statistical model estimation methods (Ljung, 1999). While there are many early papers on using machine learning to adapt parameters of continuous functions from observations (Cohen & Grossberg, 1983; Rico-Martinez & Kevrekidis, 1993; Anderson et al., 1996; González-García et al., 1998), using deep architectures to learn from sequential data has largely been dominated by Recurrent Neural Networks (RNNs) (LeCun et al., 2015) (with notable exceptions (Oord et al., 2016; Bai et al., 2018; Vaswani et al., 2017; Gu et al., 2022)). However, with the introduction of Neural Ordinary Differential Equations (neural ODEs) (Chen et al., 2018), the idea of learning continuous functions using Neural Networks (NNs) has attracted the attention of several researchers. While neural ODEs have been successfully applied in image classification (Gholami et al., 2019; Zhang et al., 2019; Zhuang et al., 2020; Gusak et al., 2020), motivated by their syntactic similarities to that of Residual Networks (ResNets) (Chen et al., 2018; Gholami et al., 2019; Zhang et al., 2019; Queiruga et al., 2020; Ott et al., 2021), their strong inductive bias towards sequential modeling makes them attractive for applications that deal with sequential data (Rubanova et al., 2019; Yildiz et al., 2019; Kidger et al., 2020; Huang et al., 2020; Chen et al., 2021; Westny et al., 2023).
@@ 1.1 Contributions
Implementations are made publicly available [^2] https://github.com/westny/neural-stability.
@@ 1.2 Related Work
Neural ODEs share several similarities with Deep State Space Models (SSMs), notably by using a continuous-time formulation. In the study by (Gu et al., 2020), an initialization technique for SSMs that enhances their ability to capture long-range dependencies was introduced—effectively integrated into the S4 model (Gu et al., 2022). However, S4 was observed to suffer from numerical instability during autoregressive generation, which was subsequently addressed in (Goel et al., 2022) by ensuring that the state matrix is constrained to be Hurwitz. While this approach shares a thematic resemblance with our emphasis on stability, a key distinction of our work is the additional focus on the numerical solvers and their stability regions—an essential aspect when solving differential equations numerically and therefore also for the training and prediction processes.
@@ 2.2 Stability Regions of Numerical Solvers
where $h$ is the step size and $\lambda$ is the eigenvalue (Ascher & Petzold, 1998, pp. 87–89). From Figure 1 it can, e.g., be deduced which maximum step size $h$ is possible given the dynamics of a system. Here, only Runge-Kutta methods are considered, but the approach can be directly extended to other integration methods.
@@ 3 Implications for Neural ODEs
Figure 2: The figure depicts the scaled poles (green crosses) of a system $f$ and the simulated response to a small perturbation using a RK4 and EF method.
The concept of stability regions is an important property of numerical ODE solvers (Ascher & Petzold, 1998). However, these stability regions and their implications on the learning and prediction performance of neural ODEs have yet to be investigated in detail. In Figure 2, the poles of a linearized nonlinear autonomous system $f$, modeled as a feedforward NN, are depicted in the complex plane, computed as the eigenvalues of $\bm{A}=\frac{\partial f(\bm{x})}{\partial\bm{x}}|_{\bm{x}=\bm{0}}$ and scaled by the step size $h$. The eigenvalues are well within the stability region of RK4 and simulating it using that solver yields a stable response. However, for the EF method, two eigenvalues are outside of its stability region, and the numerical solution exhibits oscillatory behavior resulting in large prediction errors. This is an important property to highlight in the context of neural ODEs, since training a model using a higher-order method might yield a model with eigenmodes that are problematic for a lower-order method. For example, in (Queiruga et al., 2020; Ott et al., 2021) it was observed that changing the numerical solver during inference from the solver employed during training can have a profound impact on the overall performance. While they concluded that the behavior was due to the discretization error, this could also be attributed to the solvers’ stability regions (Mohammadi et al., 2023).
@@ 3.1 Learned Model Dynamics
The choice of numerical solver affects the placement of the learned model dynamics. To illustrate, Figure 3 shows the kernel density estimate of the eigenvalues, for a linearization around a zero reference, of $400$ different neural ODEs, plotted in the complex plane. The models are parametrized using feedforward NNs and trained using the EF, Midpoint (MP), and 3rd-order Runge-Kutta (RK3) methods, respectively. They are all tasked with learning the dynamics of a linear system (one unique reference system for each model) with $3$ dynamic states that are unattainable for each solver since the eigenvalues of the state-transition jacobian of the reference systems, shown with green crosses, are manually placed outside of the respective solvers’ stability region. With some slight abuse of notation, the eigenvalues of the state-transition jacobian will be referred to as the system poles.
@@
Regardless of the nature of the system that the model should learn, the poles are contained within the stability region of the solver. This follows intuition since poles outside of the respective stability regions would yield poor prediction performance (c.f. Section 3).
@@
These findings motivate the potential need of developing initialization strategies more suitable for neural ODEs used to learn dynamic systems. Since the poles of these models are approximately contained within the stability region of the employed solver, it makes intuitive sense to initialize the model within that same region, while also adhering to the stability requirements of dynamic systems. Furthermore, since the model is stable from the start of training, this allows for training with longer prediction horizons without risking divergence. In addition, motivated by the increased difficulty to learn fast dynamics, we hypothesize that by initializing the model with poles further into the left-half plane the model will experience a smoother learning process.
@@ 4.1 Parameter Initialization
which is exactly an $n$-layer linear feedforward network. This expression is central to the initialization technique—outlined next. Without loss of generality and for simplicity of the presentation, $\kappa$ is from now on assumed 1.
@@
Then, clearly, $\bm{A}$ in (5) has the desired eigenvalues $\lambda_{j}$. The $n$th root is used to allocate information about each eigenvalue throughout the whole feedforward network. This constitutes one alternative to initializing the weights $\bm{W}_{i}$ but the approach has several restrictions: 1) the desired eigenvalues may not be real, 2) the hidden layer dimensions may not all be the same, and 3) it is undesired that the initialized weight matrices have all off-diagonal elements set to 0.
First, consider the case with non-real eigenvalues. With real weights $\bm{W}_{i}\in\mathbb{R}^{{d_{h_{i}}\times d_{h_{i-1}}}}$ in (4) it is required that complex eigenvalues come in conjugate pairs. Consequently, when $d_{x}$ is odd, the system must have a least one real eigenvalue. Regardless, the construction is still straightforward. Let $\lambda_{k}\in\mathbb{C}$ and, as before, compute the $n$th (principal) root as $\lambda_{k}^{1/n}=\mu_{k}+j\omega_{k}$ where $j$ is the imaginary unit. First, note that the matrix
@@
Now, $\text{dim}(\bm{W}_{i})=\text{dim}(\bm{A})$ is rarely the case when constructing a feedforward network like (4). Secondly, to avoid sparse weight matrices where there is a risk of slow learning due to the vanishing gradient problem (Basodi et al., 2020; Pascanu et al., 2013), we would like to avoid matrix elements with a value of zero. Finally, similar to established initialization techniques (Glorot & Bengio, 2010; Saxe et al., 2013; He et al., 2015), adding randomness to the weight initialization assures that diverse models can be generated, but also crucially breaks symmetry among neurons—enabling diverse feature extraction. To meet these requirements, the previous ideas presented above are generalized as
$\bm{A}=\underbrace{\Lambda_{n}\Pi_{n-1}^{-1}}_{\bm{W}_{n}}\underbrace{\Pi_{n-1}\Lambda_{n-1}\Pi_{n-2}^{-1}}_{\bm{W}_{n-1}}\cdots\underbrace{\Pi_{2}\Lambda_{2}\Pi_{1}^{-1}}_{\bm{W}_{2}}\underbrace{\Pi_{1}\Lambda_{1}}_{\bm{W}_{1}}$ (9)
where $\Lambda_{i}$ is used to address the dimension property and $\Pi_{i}$ the sparseness (randomness) property.
The matrices $\Lambda_{i}\in\mathbb{R}^{{d_{h_{i}}\times d_{h_{i-1}}}}$ are constructed to inherit the properties of (6) and (8) for real and complex eigenvalues respectively such that they contain a factor of the system eigenvalues. To ensure that the dimensions of the respective matrix products adhere to the hidden dimensions of the layers in the network, the matrices are extended with blocks of zeros. Consider for example the simple case that $d_{x}=2$ and $d_{h_{1}}=4$, then
@@
Let the matrices $\Pi_{i}\sim\mathcal{O}(d_{h_{i}})$ where $\mathcal{O}$ is the Haar distribution of orthogonal matrices (Mezzadri, 2006). The choice of using orthogonal matrices is natural since their eigenvalues being $1$ ensures that the complete model’s eigenvalues are neither attenuated nor amplified along the computational graph and their property of having their transpose equal to their inverse offers numerical and implementational benefits. Note that introducing the random matrices $\Pi_{i}$ does not change (9) since
@@
although their inclusion assures that $\forall\text{w}_{j,k}\in\bm{W}_{i},\text{w}_{j,k}\neq 0$ with probability 1.
@@
To include external inputs the vector $\bm{u}$ is assumed to be concatenated with the state vector $\bm{x}$ prior to propagation. This is easily accounted for by modifying $\Lambda_{1}$ such that $\Lambda_{1}\in\mathbb{R}^{{d_{h_{1}}\times(d_{x}+d_{u})}}$. Continuing the example in (10), with $d_{u}=1$, then
@@
[…]
@@ 5 Teacher-Student Regression
We examine a teacher-student NN structure in which a student network is trained to replicate the output of a teacher network designed to reflect a nonlinear dynamic system. This system is adaptable to various dynamic states and inputs, allowing for a rich simulation study. Both the teacher and student models are outfitted with nonlinear ELU (Clevert et al., 2015) activation functions, which are interposed within the network layers. While different setups were investigated, the main considered input was in the form of pulse-width modulated waveforms $w$, meant to model a continuous-time input that excites model dynamics. The training and test set inputs $\bm{u}$ are constructed using different periodic functions with varying frequencies together with added noise
@@
where $v\sim\mathcal{N}(0,0.1)$ and $N$ is the number of samples. Test data generation then proceeds by solving the Initial Value Problem (IVP)
@@ Results
In general, the eigenvalues of the true system to be learned from training data are rarely known. To simulate the case when neither the solver nor the step size is chosen optimally, the second investigation is conducted to study how well the nonlinear students can approximate the data regardless. This is done by initializing the teacher model with $2$ dynamic states using the proposed technique such that the poles are outside of the stability region of the employed solver. To add to the difficulty of the task, each student model is under-parameterized in comparison with the teacher. While the teacher has $3$ layers with $128$ nodes in each layer, the student has $2$ layers with $32$ nodes in each layer.
@@ 6 Experiments
To assess the performance and usability of the proposed technique, several experiments are conducted across a diverse set of problem domains. For each experimental scenario, multiple instances of the models are trained under different random seeds to ensure the reproducibility and statistical significance of the results. Importantly, the configuration of hyperparameters, e.g. batch size and learning rate may differ with the task but is always the same for all methods. Unless stated otherwise, the models are trained and tested using the Midpoint method.

### v3

@@ 1 Introduction
Learning dynamic systems from observations has a long history of applications, particularly in time-series analysis and prediction. This research domain is intrinsically connected to classical system identification and statistical model estimation methods (Ljung, 1999). While there are many early studies on using machine learning to adapt parameters of continuous functions from observations (Cohen & Grossberg, 1983; Rico-Martinez & Kevrekidis, 1993; Anderson et al., 1996; González-García et al., 1998), using deep architectures to learn from sequential data has largely been dominated by Recurrent Neural Networks (RNNs) (LeCun et al., 2015), with notable exceptions (Oord et al., 2016; Bai et al., 2018; Vaswani et al., 2017; Gu et al., 2022). However, with the introduction of Neural Ordinary Differential Equations (neural ODEs) (Chen et al., 2018), the idea of learning continuous functions using Neural Networks (NNs) has attracted the attention of several researchers. While neural ODEs have been successfully applied in image classification (Gholami et al., 2019; Zhang et al., 2019; Zhuang et al., 2020; Gusak et al., 2020), motivated by their syntactic similarities to that of Residual Networks (ResNets) (Chen et al., 2018; Gholami et al., 2019; Zhang et al., 2019; Queiruga et al., 2020; Ott et al., 2021), their strong inductive bias towards sequential modeling makes them attractive for applications that deal with sequential data (Rubanova et al., 2019; Yildiz et al., 2019; Kidger et al., 2020; Huang et al., 2020; Chen et al., 2021; Westny et al., 2023; Lipman et al., 2023; Verma et al., 2024).
@@ 1.1 Contributions
Implementations are made publicly available. [^2] https://github.com/westny/neural-stability
@@ 1.2 Related Work
Neural ODEs share several similarities with Deep State Space Models (SSMs), notably by using a continuous-time formulation. In the study by (Gu et al., 2020), an initialization technique for SSMs that enhances their ability to capture long-range dependencies was introduced—effectively integrated into the S4 model (Gu et al., 2022). However, S4 was observed to suffer from numerical instability during autoregressive generation, which was subsequently addressed in (Goel et al., 2022) by ensuring that the state matrix is constrained to be Hurwitz. Similarly, one can also draw connections between this work and methods in the reservoir computing literature (Lukoševičius & Jaeger, 2009) as well as the choice of stable dynamical systems in Random Projection Filter Banks (Farahmand et al., 2017). While these approaches share a thematic resemblance with our emphasis on stability, a key distinction of our work is the additional focus on the numerical solvers and their stability regions—an essential aspect when solving differential equations numerically and therefore also for the training and prediction processes.
@@ 2.2 Stability Regions of Numerical Solvers
where $h$ is the step size and $\lambda$ is the eigenvalue (Ascher & Petzold, 1998, pp. 87–89). From Figure 1 it can, e.g., be deduced which maximum step size $h$ is possible given the dynamics of a system. Here, only fixed-step explicit methods are considered, but the approach can be directly extended to other integration methods. For example, several implicit solvers have well-defined stability regions (Ascher & Petzold, 1998, p. 143).
@@ 3 Implications for Neural ODEs
Figure 2: The figure depicts the scaled poles (green crosses) of a synthetic dynamic system $f$ and the simulated response to a small perturbation using a RK4 and EF method.
The concept of stability regions is an important property of numerical ODE solvers (Ascher & Petzold, 1998). However, these stability regions and their implications on the learning and prediction performance of neural ODEs have yet to be investigated in detail.
Consider a dynamic system $\dot{\bm{x}}=f(\bm{x})$, modeled using a feedforward NN with nonlinear activations. In Figure 2, we illustrate the poles of the linearized system within the complex plane, i.e., the eigenvalues of the Jacobian matrix $\bm{A}=\frac{\partial f(\bm{x})}{\partial\bm{x}}|_{\bm{x}=\bm{0}}$, scaled by the step size $h$. Clearly, the eigenvalues are well within the stability region of RK4 and simulating it using that particular solver yields a stable response. However, for the EF method, the two complex eigenvalues are outside of its stability region, and the simulated response exhibits oscillatory behavior resulting in large prediction errors. This is an important property to highlight in the context of neural ODEs, since training a model using a higher-order method might yield a model with eigenmodes that are problematic for a lower-order method. For example, in (Queiruga et al., 2020; Ott et al., 2021) it was observed that changing the numerical solver during inference from the solver employed during training can have a profound impact on the overall performance. While they concluded that the behavior was due to the discretization error, this could also be attributed to the solvers’ stability regions (Mohammadi et al., 2023).
@@ 3.1 Learned Model Dynamics
The choice of numerical solver affects the placement of the learned model dynamics. To illustrate, Figure 3 shows the kernel density estimate of the eigenvalues, for a linearization around a zero reference of $400$ different neural ODEs, depicted in the complex plane. The models are parametrized using feedforward NNs and trained using the EF, Midpoint (MP), and 3rd-order Runge-Kutta (RK3) methods, respectively. They are all tasked with learning the dynamics of a linear system (one unique reference system for each model) with $3$ dynamic states that are unattainable for each solver since the eigenvalues of the state-transition jacobian of the reference systems, shown with green crosses, are manually placed outside of the respective solvers’ stability region. With some slight abuse of notation, the eigenvalues of the state-transition jacobian will be referred to as the system poles.
@@
Regardless of the nature of the system that the model should learn, the poles are contained within the stability region of the solver. This follows intuition since poles outside of the respective stability regions would yield poor prediction performance (cf. Section 3).
@@
These findings motivate the potential need to develop initialization strategies more suitable for neural ODEs used to learn dynamic systems. Since the poles of these models are approximately contained within the stability region of the employed solver, it makes intuitive sense to initialize the model within that same region, while also adhering to the stability requirements of dynamic systems. Furthermore, since the model is stable from the start of training, this allows for training with longer prediction horizons without risking divergence. In addition, motivated by the increased difficulty of learning fast dynamics, we hypothesize that by initializing the model with poles further into the left-half plane the model will experience a smoother learning process.
@@ 4.1 Parameter Initialization
Note that the stability of the linearized system does not necessarily imply the stability of the nonlinear system, or even of a linear time-varying system, as discussed in (Khalil, 2002). However, analyzing the stability of the linearized system can provide a useful starting point for initialization.
@@
which is exactly an $n$-layer linear feedforward network. This expression is central to the initialization technique—outlined next. Without loss of generality and for simplicity of the presentation, $\kappa$ is from now on assumed 1. Note that, $\bm{W}_{i}$ should be scaled by $1/\kappa_{i}$ if $\kappa_{i}\neq 1$.
@@
Then, clearly, $\bm{A}$ in (5) has the desired eigenvalues $\lambda_{j}$. The $n$th root is used to allocate information about each eigenvalue throughout the whole feedforward network. This constitutes one alternative to initializing the weights $\bm{W}_{i}$ but the approach has several restrictions, described next.
First, consider the case when one wants the eigenvalues to be complex. With real weights $\bm{W}_{i}\in\mathbb{R}^{{d_{h_{i}}\times d_{h_{i-1}}}}$ in (4), this requires that complex eigenvalues come in conjugate pairs. Consequently, when $d_{x}$ is odd, the system must have a least one real eigenvalue. Regardless, the construction is still straightforward. Let $\lambda_{k}\in\mathbb{C}$ and, as before, compute the $n$th (principal) root as $\lambda_{k}^{1/n}=\mu_{k}+j\omega_{k}$ where $j$ is the imaginary unit. First, note that the matrix
@@
Next, we need to address the dimensionality and sparsity of the weight matrices. Certainly, $\text{dim}(\bm{W}_{i})\neq\text{dim}(\bm{A})$ is most often the case when constructing a feedforward network like (4). Additionally, to avoid sparse weight matrices where there is a risk of slow learning due to the vanishing gradient problem (Basodi et al., 2020; Pascanu et al., 2013), we would like to avoid matrix elements with a value of zero. Finally, similar to established initialization techniques (Glorot & Bengio, 2010; Saxe et al., 2014; He et al., 2015), adding randomness to the weight initialization assures that diverse models can be generated, but also crucially breaks symmetry among neurons—enabling diverse feature extraction. To meet these requirements, the previous ideas presented above are generalized as
$\bm{A}=\underbrace{\Lambda_{n}\Pi_{n-1}^{-1}}_{\bm{W}_{n}}\underbrace{\Pi_{n-1}\Lambda_{n-1}\Pi_{n-2}^{-1}}_{\bm{W}_{n-1}}\cdots\underbrace{\Pi_{2}\Lambda_{2}\Pi_{1}^{-1}}_{\bm{W}_{2}}\underbrace{\Pi_{1}\Lambda_{1}}_{\bm{W}_{1}},$ (9)
where $\Lambda_{i}$ is used to address the dimension property and $\Pi_{i}$ the denseness and randomness property.
The matrices $\Lambda_{i}\in\mathbb{R}^{{d_{h_{i}}\times d_{h_{i-1}}}}$ are constructed to inherit the properties of (6) and (8) for real and complex eigenvalues respectively such that they contain a factor of the system eigenvalues. To ensure that the dimensions of the respective matrix products adhere to the hidden dimensions of the layers in the network, the matrices are extended with blocks of zeros. Consider for example the simple case when $d_{x}=2$ and $d_{h_{1}}=4$, then
@@
The matrices $\Pi_{i}\in\mathbb{R}^{{d_{h_{i}}\times d_{h_{i}}}}$ are random orthogonal matrices sampled from the Haar distribution (Stewart, 1980; Mezzadri, 2006). In practice, these are constructed by applying QR decomposition of a $d_{h_{i}}\times d_{h_{i}}$ matrix whose elements are independently normally distributed with zero mean and unit variance (Mezzadri, 2006). The choice of using orthogonal matrices is natural since their eigenvalues being $1$ ensures that the complete model’s eigenvalues are neither attenuated nor amplified along the computational graph and their property of having their transpose equal to their inverse offers numerical and implementational benefits. Note that introducing the random matrices $\Pi_{i}$ does not change (9) since
@@
although their inclusion assures that $\forall\text{w}_{j,k}\in\bm{W}_{i},\text{w}_{j,k}\neq 0$ with probability 1. Although not explored in this work, we note that it is possible to use other eigenvectors than the standard basis when constructing $\bm{A}$.
@@
To include external inputs, the vector $\bm{u}$ is assumed to be concatenated with the state vector $\bm{x}$ prior to propagation. This is easily accounted for by modifying $\Lambda_{1}$ such that $\Lambda_{1}\in\mathbb{R}^{{d_{h_{1}}\times(d_{x}+d_{u})}}$. Continuing the example in (10), with $d_{u}=1$, then
@@
Now that a method to randomly initialize the feedforward network (4) given a set of eigenvalues has been outlined, the procedure is completed using a simple rejection sampling algorithm. This involves randomizing real and/or complex-conjugate eigenvalue pairs within the stability region of the selected solver for a given step size based on condition (3). This procedure is outlined in Algorithm 1. The rejection-sampling algorithm produces a set of eigenvalues that are utilized to form $\Lambda_{i}$ as described in (10) and (12). Subsequently, random orthogonal matrices $\Pi_{i}$ are sampled. These matrices, together with $\Lambda_{i}$ are used to construct the weights $W_{i}$ as outlined in (9). The process concludes with the construction of the bias terms as in (13).
Algorithm 1 Stability Region Rejection Sampling
  Input: Solver order $p$, step size $h$, state dimension $d_{x}$
  Initialize: ${}\mathcal{E}\leftarrow\{\}$
  repeat
     Sample: $\mu\sim\mathcal{U}(-3.0,-0.1)$
     if $|\mathcal{E}|<d_{x}-2$ and use_complex then
        Sample: $\omega\sim\mathcal{U}(-3.0,3.0)$
     else
        Set: $\omega=0$
     end if
     Form: $\lambda=\mu+j\omega$
     if $\left|1+\lambda+\frac{\lambda^{2}}{2}+\cdots+\frac{\lambda^{p}}{p!}\right|<1-\epsilon$ then
        $\mathcal{E}\leftarrow\mathcal{E}\cup\{\lambda/h\}$
        if $\omega\neq 0$ then
           $\mathcal{E}\leftarrow\mathcal{E}\cup\{\bar{\lambda}/h\}$
        end if
     else
        Reject and proceed to the next iteration
     end if
  until $|\mathcal{E}|=d_{x}$
  Return: $\mathcal{E}$
@@ 5 Teacher-Student Regression
We examine a teacher-student NN setting in which a student network is trained to replicate the output of a teacher network. The teacher is designed to simulate a nonlinear dynamic system and is adaptable to various dynamic states and inputs, allowing for a rich simulation study. Both the teacher and student models are outfitted with nonlinear ELU (Clevert et al., 2015) activation functions. The main considered input was in the form of pulse-width modulated waveforms $w$, meant to model a continuous-time input that excites model dynamics. The training and test set inputs $\bm{u}$ are constructed using different periodic functions with varying frequencies together with added noise
@@
where $v\sim\mathcal{N}(0,0.1)$ and $N$ is the number of samples. Data generation then proceeds by solving the Initial Value Problem (IVP)
@@ Results
In general, the eigenvalues of the true system to be learned from training data are rarely known. To simulate the case when neither the solver nor the step size is chosen optimally, the second investigation is conducted to study how well the nonlinear students can approximate the data regardless. This is done by initializing the teacher model with $2$ dynamic states using the proposed technique such that the poles are outside of the stability region of the employed solver. To add to the difficulty of the task, each student model is under-parameterized relative to the teacher. While the teacher has $3$ layers with $128$ nodes in each layer, the student has $2$ layers with $32$ nodes in each layer.
@@ 6 Experiments
To assess the performance and usability of the proposed technique, several experiments are conducted across a diverse set of problem domains. Our choice of experimental cases was guided by three principal criteria: (i) prevalence in the literature, (ii) complexity and diversity of the tasks, and (iii) real-world applicability and relevance. For each experimental scenario, multiple instances of the models are trained under different random seeds to ensure the reproducibility and statistical significance of the results. Importantly, the configuration of hyperparameters, e.g., batch size and learning rate may differ with the task but is always the same for all methods. Unless stated otherwise, the models are trained and tested using the Midpoint method.